# Drift No More? Context Equilibria in Multi-Turn LLM Interactions

### v1

@@ Abstract
Drift No More? Context Equilibria in Multi-Turn LLM Interactions
Vardhan Dongre1,2 Ryan A. Rossi2 Viet Dac Lai2 David Seunghyun Yoon2 Dilek Hakkani-Tür1 Trung Bui2 1University of Illinois Urbana-Champaign 2Adobe Research
@@ 1 Introduction
Large Language Models (LLMs) have become central to a wide range of interactive systems, from virtual assistants and copilots to autonomous agents Ouyang et al. (2022); Achiam et al. (2023); Brown et al. (2020) that plan Yao et al. (2023); Wang et al. (2023c); Li et al. (2023); Dongre et al. (2024), explain Cai et al. (2019), or negotiate Lewis et al. (2017); Bianchi et al. (2024) over extended dialogues. Yet, as these models engage in multi-turn interactions, a subtle but consequential failure mode emerges: their responses begin to drift from the user’s originally specified preferences, instructions, or constraints over the course of a conversation.
Unlike factual hallucinations Ji et al. (2023) or local coherence errors, context drift is a slow erosion of intent: a summarizer that gradually loses the requested tone, an image editing agent that drifts from the target aesthetic in an image, and a user simulator that forgets its goals and behavioral constraints. Critically, most current benchmarks and evaluations are blind to this degradation, focusing either on end-task success Thoppilan et al. (2022); Zhou et al. (2023) or per-turn quality Guan et al. (2025); Kwan et al. (2024); Dongre et al. (2025); Wang et al. (2023b); Chang et al. (2024); Duan et al. (2023), without capturing temporal misalignment across turns.
The prevailing intuition is that context drift accumulates unboundedly as conversations lengthen, owing to memory limits, information loss, or compounding errors. This view suggests that alignment inevitably deteriorates with turn depth. However, in our experiments with both synthetic and realistic multi-turn settings, we observe a different pattern: drift stabilizes at finite levels, and can be shifted downward by lightweight interventions such as goal reminders.
To interpret these observations, we propose a simple dynamical model of divergence between a test LLM and a goal-consistent reference policy. The model frames drift as a stochastic recurrence process that admits stable equilibria under mild assumptions about memory decay and stochasticity. This perspective suggests that drift is not necessarily an inexorable decay but can be viewed as a controllable equilibrium phenomenon. Our contributions in this work can be summarized as follows:
@@ 2 Related Works
A growing body of work has identified that large language models can suffer performance loss in extended conversations. Laban et al. (2025) show that model outputs gradually deviate from earlier context, often leading to incoherence or goal neglect. Abdelnabi et al. (2024) measure “task drift” by tracking changes in model activations over turns and propose detection mechanisms to flag when models are likely to have lost the original task. Mehri et al. (2025) examine goal consistency over long-horizon dialogue with user simulators and call it "instruction drift", highlighting that even capable models struggle to sustain alignment as conversations deepen. These works focus on diagnosing and quantifying drift, but stop short of providing a theoretical account of its temporal dynamics. In contrast, we propose a simple dynamical perspective that models drift as a bounded stochastic process rather than as inevitable monotonic decay. Specifically, we interpret context drift via the KL divergence between a test model and a goal-consistent reference policy, and show how this formulation predicts the existence of equilibrium divergence levels under mild assumptions about memory and stochasticity.
@@
Recent studies Zhang and Dong (2024); Bhargava et al. (2023); Li et al. (2024a) have aimed to formalize LLM behavior through the lens of dynamical systems and control theory. Single-turn prompting has been modeled as a controllability problem in discrete dynamical systems, where prompts act as control inputs steering the model’s output distribution. Bhargava et al. (2023) treat transformer-based LLMs as discrete stochastic dynamical systems and analyze the controllability of self-attention, showing how short prompts can dramatically steer reachable outputs. Zhang and Dong (2024) extend this perspective by modeling transformers via Neural ODEs and integrating robust control methods to stabilize outputs. Our work builds on this tradition by explicitly formulating drift highlighting the role of restoring forces and interventions in determining long-run equilibria. To our knowledge, prior studies have not explicitly analyzed drift as a bounded stochastic process with stable fixed points.
@@
Another strand of work attributes multi-turn failures to imperfect memory mechanisms. Studies on memory-augmented models Wang et al. (2023a); Li et al. (2024b) and context compression Jiang et al. (2023a, b) investigate ways of preserving salient information. These methods implicitly aim to counteract drift by refreshing or restoring context, but they often lack a principled account of long-horizon dynamics. Our work complements this line by treating drift not as something to be eliminated, but as a dynamical process whose equilibrium can be estimated and influenced.
@@ 6 Results
| Model | KL Divergence | JS Divergence | Sim | LLM Judge Score |
@@ 10 Future Work
Appendix A Appendix
A.1 Proof Sketch of Bound
We sketch the reasoning behind Eq. 3. Under Eq. 1, assuming is monotone and , we can write
for some contraction factor . Unrolling this recursion over steps yields
which gives the stated inequality. The result is illustrative rather than universal: it shows that bounded noise leads to convergence to a finite equilibrium, and that positive interventions shift the equilibrium downward.
A.2 Linear Drift Diagnostic
Starting from the recurrence model in Eq. (1):
we linearize around the equilibrium :
Substituting and taking expectations under bounded noise gives:
Grouping constants yields the empirical form
where and . The empirical equilibrium thus estimates the fixed point where .
Appendix B Statistical Reliability of Fitted Coefficients
For each model and condition, we estimate via ordinary least squares (OLS) and compute 95% confidence intervals using bootstrapping over conversation trajectories. Across all settings, the sign of remains consistently negative within the confidence bounds, indicating robustness of the restoring-force interpretation. Average values range from 0.28–0.72 (Table 6), showing that the linear model captures a substantial fraction of variance in given the stochasticity of generation.
Appendix C Tasks
C.1 Synthetic constrained multi-turn generation task
The synthetic task is designed to let us precisely observe and manipulate drift in a controlled environment, where the ground truth goal is unambiguous and drift can be induced in a measurable way. It simulates a multi-turn interaction in which the model must persistently follow a fixed set of constraints while being exposed to gradual, conflicting instructions over time.
Turn-wise Behavior and Interventions:
Table 5 shows a trajectory comparing GPT-4.1 (reference) and LLaMA-3.1-8B (test) across four turns. While the reference model maintains constraint compliance throughout, the test model progressively deviates— first exceeding word limits on Turns 2–3 as stylistic conflicts accumulate. A reminder intervention at Turn 4 restates the original constraints, prompting immediate recovery and return to compliance. This pattern demonstrates the key dynamics predicted by our framework: drift arises gradually through compounding contextual pressures but can be corrected by minimal, well-timed interventions ().
| Synthetic Task Example Original passage. “Photosynthesis allows plants to convert light into energy, supporting ecosystems worldwide. Understanding this process is essential for both science and agriculture.” Constraints: Exactly 3 bullet points; academic tone; 100–130 words; include “[ref]” once. A reminder restating constraints is injected at Turn 4. | ||
| Turn | GPT-4.1 (Reference) | LLaMA-3.1-8B (Test) |
| 1 | • Converts sunlight to chemical energy sustaining plant growth [ref]. • Enables ecosystem stability and supports food systems. • Study informs sustainable practice in agriculture. ( 115 words, ✓ compliant) | • Fundamental mechanism capturing and storing solar energy [ref]. • Supports trophic chains and ecological resilience. • Core topic in agronomic and climate research. ( 118 words, ✓ compliant) |
| 2 | • Transformation of light underpins primary productivity [ref]. • Reliability depends on stable long-horizon behavior. • Research connects mechanisms to field outcomes. ( 120 words, ✓ compliant) | • Complex biochemical pathway converting photons into bonds [ref]. • Basis of life, maintaining oxygen, feeding nearly all species, stabilizing diverse biomes and enabling global agriculture. • Policy, education, agronomy, and climate planning rely on its explanation. ( 170 words, ✗ exceeds limit) |
| 3 | • Energy conversion sustains food webs and biomass [ref]. • Regulates atmospheric balance and local climates. • Efficiency matters for yields and sustainability. ( 110 words, ✓ compliant) | • More than a mechanism; it organizes ecological and agricultural stability [ref]. • Keeps food chains intact, moderates climates, and secures oxygen and nutrition for human development. • Understanding dynamics informs coordinated sustainability goals. ( 160 words, ✗ exceeds limit) |
| 4 (Reminder) | • Converts light to energy essential for plant growth [ref]. • Maintains ecological balance and food security. • Ongoing study advances sustainable agriculture. ( 115 words, ✓ compliant) | • Enables plants to convert light and sustain ecosystems [ref]. • Reminder restores concise academic style and structure. • Output returns to the specified three-bullet format. ( 125 words, ✓ back in range) |
| Model | Condition | Residual Std. | Max Residual | Spearman | ||||
| GPT-4.1 | Baseline | |||||||
| GPT-4.1 | Reminders | |||||||
| Llama-3.1-70B | Baseline | |||||||
| Llama-3.1-70B | Reminders | |||||||
| Llama-3.1-8B | Baseline | |||||||
| Llama-3.1-8B | Reminders |
C.2 -Bench Setup
We leverage -Bench Yao et al. (2024) as a benchmark framework for realistic goal-driven dialogues in structured domains such as retail order management and airline reservations. -Bench provides (i) task-oriented agents with tool APIs (e.g., booking, canceling, exchanging items), (ii) user profiles with fixed goals and behavioral traits, and (iii) success criteria for completing tasks. See Figure 6 for further details.
Simulation Protocol.
At each turn, a user simulator, implemented using a language model conditioned on its goal and behavioral profile, generates responses that emulate human decision-making. The tool-using agent interacts with this simulator through -Bench APIs (e.g., booking, checking availability, or processing exchanges). The reference policy, instantiated with GPT-4.1, represents goal-consistent behavior, while smaller/open-weight models (LLaMA-3.1-8B, LLaMA-3.1-70B, Qwen-2-7B-Instruct) serve as test simulators. Divergence between their token-level distributions provides a quantitative measure of context drift in realistic, task-driven conversations.
Metrics and Interventions.
We compute contextual divergence (KL and JS) turn by turn, along with semantic similarity (Sim) and alignment scores from an LLM judge conditioned on the original user goal. To test drift controllability, explicit goal-reminder interventions are injected at fixed turns ( and ). Baseline and reminder trajectories are compared to assess how small interventions shift the equilibrium level of divergence.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have become central to a wide range of interactive systems, from virtual assistants and copilots to autonomous agents (Ouyang et al. 2022; Achiam et al. 2023; Brown et al. 2020; Acikgoz et al. 2025) that plan (Yao et al. 2023; Wang et al. 2023c; Li et al. 2023; Dongre et al. 2024), explain (Cai et al. 2019), or negotiate (Lewis et al. 2017; Bianchi et al. 2024) over extended dialogues. Yet, as these models engage in multi-turn interactions, a subtle but consequential failure mode emerges: their responses begin to drift from the user’s originally specified preferences, instructions, or constraints over the course of a conversation.
Unlike factual hallucinations (Ji et al. 2023) or local coherence errors, context drift is a slow erosion of intent: a summarizer that gradually loses the requested tone, an image editing agent that drifts from the target aesthetic in an image, and a user simulator that forgets its goals and behavioral constraints. Critically, most current benchmarks and evaluations are blind to this degradation, focusing either on end-task success (Thoppilan et al. 2022; Zhou et al. 2023) or per-turn quality (Guan et al. 2025; Kwan et al. 2024; Dongre et al. 2025; Wang et al. 2023b; Chang et al. 2024; Duan et al. 2023), without capturing temporal misalignment across turns.
The prevailing intuition is that context drift accumulates unboundedly as conversations lengthen, owing to memory limits, information loss, or compounding errors. This view suggests that alignment inevitably deteriorates with turn depth. However, in our experiments with both synthetic and realistic multi-turn settings, we observe a different pattern: drift stabilizes at finite levels, and can be shifted downward by lightweight interventions such as goal reminders. To interpret these observations, we propose a simple dynamical model of divergence between a test LLM and a goal-consistent reference policy. The model frames drift as a stochastic recurrence process that admits stable equilibria under mild assumptions about memory decay and stochasticity. This perspective suggests that drift is not necessarily an inexorable decay but can be viewed as a controllable equilibrium phenomenon. Our contributions in this work can be summarized as follows:
@@ 2 Related Works
A growing body of work has identified that large language models can suffer performance loss in extended conversations. (Laban et al. 2025) show that model outputs gradually deviate from earlier context, often leading to incoherence or goal neglect. (Abdelnabi et al. 2024) measure “task drift” by tracking changes in model activations over turns and propose detection mechanisms to flag when models are likely to have lost the original task. (Mehri et al. 2025) examine goal consistency over long-horizon dialogue with user simulators and call it "instruction drift", highlighting that even capable models struggle to sustain alignment as conversations deepen. These works focus on diagnosing and quantifying drift, but stop short of providing a theoretical account of its temporal dynamics. In contrast, we propose a simple dynamical perspective that models drift as a bounded stochastic process rather than as inevitable monotonic decay. Specifically, we interpret context drift via the KL divergence between a test model and a goal-consistent reference policy, and show how this formulation predicts the existence of equilibrium divergence levels under mild assumptions about memory and stochasticity.
@@
Recent studies (Zhang and Dong 2024; Bhargava et al. 2023; Li et al. 2024a) have aimed to formalize LLM behavior through the lens of dynamical systems and control theory. Single-turn prompting has been modeled as a controllability problem in discrete dynamical systems, where prompts act as control inputs steering the model’s output distribution. (Bhargava et al. 2023) treat transformer-based LLMs as discrete stochastic dynamical systems and analyze the controllability of self-attention, showing how short prompts can dramatically steer reachable outputs. (Zhang and Dong 2024) extend this perspective by modeling transformers via Neural ODEs and integrating robust control methods to stabilize outputs. Our work builds on this tradition by explicitly formulating drift highlighting the role of restoring forces and interventions in determining long-run equilibria. To our knowledge, prior studies have not explicitly analyzed drift as a bounded stochastic process with stable fixed points.
@@
Another strand of work attributes multi-turn failures to imperfect memory mechanisms. Studies on memory-augmented models (Wang et al. 2023a; Li et al. 2024b) and context compression (Jiang et al. 2023a, b) investigate ways of preserving salient information. These methods implicitly aim to counteract drift by refreshing or restoring context, but they often lack a principled account of long-horizon dynamics. Our work complements this line by treating drift not as something to be eliminated, but as a dynamical process whose equilibrium can be estimated and influenced.
@@ 6 Results
| Model | KL Divergence | JS Divergence | Sim | Judge Score |